FEL Monthly Highlights
March 2001

For the last week of the winter FEL user run which ended on March 9, the FEL produced new power records for femtosecond class green light (56 watts) and blue (350nm) light (16 watts).  This light was used for patterning experiments on photosensitive glass by our user from the Aerospace Corp.  We also obtained very nice angular distributions of the Compton x-ray source.  Please check out the new and improved web site http://laser.jlab.org.  There are images and movies from the last run in the "media gallery" located on the top of the index page.

The FEL project had the pleasure of hosting an important visit on March 22:  Vice Admiral G. P. Nanos, Rear Admiral M. Mathis from the Naval Sea Systems Command, and Brig. General R. Natonski (USMC).  The group was impressed with the FEL project to date and was particularly interested in the capabilities of the existing FEL user facility for laser-materials and propagation studies of interest to the Navy.  We were told to expect a return visit and we look forward to that opportunity.

The FEL team successfully defended our progress on the design, analysis and construction activities for the 10 kW Upgrade project at the third semiannual project review held on March 27.  As a useful follow-up activity a 1-day workshop was held on March 28 on the subject of FEL power scaling activities that took advantage of the review committee's presence.

A rebaseline of the Phase 1 cost performance plan and the initial plan for the Phase 2 project was forwarded to the ONR contract monitor for review and acceptance.  The Phase 1 was rebaselined to reflect  Engineering Change Requests approved at the October 2000 project review:  (1) the descoping of the optical cavity from an R5 configuration to a simpler and less costly concentric cavity; (2) the evolution from the driver accelerator design from Release 1.0 to Release 1.1; and the (3) the change in vacuum hardware from 2 inch to 3 inch aperture.  The initial Phase 2 cost performance plan has been drafted to optimize the schedule and minimize the cost of the complete Phase 1 and Phase 2 efforts.

According to the semiannual project review committee chaired by John Albertine, “there are no problems nor surprises" in our performance to date.  The committee noted its concern for missing budget elements that are presently not in the budget scope for FY02: (1) approx. $750k to finish the 3rd FEL Upgrade cryomodule and commissioning funds. The committee also noted some concern for the design power loading levels (both peak and average) for the resonator optics.  A follow-up chart is being prepared for the committee that shows the upgrade design values are at or below power loading levels that have been already demonstrated in the 1 kW demo.  With the completion of the semiannual review, the project monthly report was completed for the month of February, which includes a rebaselined cost-performance report that was shown and approved at the review.  The rebaseline incorporates three engineering change request and the inclusion of the cost-performance plan for the Phase 2 of the project, which extends from Feb.1, 2001 to Sept. 30, 2002.  We were pleased to have John Eric from the Air Force Research Lab attend the review.  John will be the contract monitor for the UV FEL portion of the Upgrade project in April which we anticipate will be starting up sometime in April.
On Friday, March 9 we had a visit from a team from AES who presented a summary of their SBIR effort funded by SMDC on the design of a 100 mA injector based on the current 5 mA FEL injector.  We were visited on Wednesday, March 14 by a 6 member technical team from Raytheon who were interested in proposing a plan for laser development for the Navy.
Project Cost Performance:
The project (Phase 1) budget for the period June 1, 2000 to Sept. 30, 2001 is $9,029k.  The Phase 2 project was approved for $4,500 k for a performance period of Feb. 1, 2001 to Sept. 30, 2002.  The project through the month of March has a total of $2,232k of performance scheduled (assuming the project started at the originally planned start date of April 1, 2000).  The work performed through the end of March was $3,239k, which is 35% complete vs. 25% scheduled.  The actual cost accrued through March totals $4,044k.  This results in a schedule variance of +$1,006k and a cost variance of -$806k.  The growth in negative cost variance is largely due to the accumulation of design labor charges that can not be booked against earned value until subsystems are completed or delivered.

WBS 3 (Beam Physics):
A “final" solution for the Rev 1.1 UV compatible design was established, completing the initial accelerator physics design phase for the machine given our present resources.  Activities now turn to documentation, performance analysis/component specification, machine modeling and disaster management “What if CSR DOES matter?").  Machine studies time was used to resurrect difference orbit measurements and apply them to a carefully established machine diagnostic configuration.

Paperwork was generated for the Rev 1.1 ECR.  Discussions were held about diagnostic systems layout. Archival DIMAD design files for the IR (032101/032101out) and UV (uv0313/uv313out).

Injection line dipole specifications were analyzed and will be documented in a forthcoming technical note.
We also started performance analysis on the IR Rev 1.1 design, which led immediately to work on optimization of the sextupole correction package.  Further work was performed on beam loss instability analysis.

WBS 4 (Injector):
Ordered flanges and ion pumps for the gun chamber.  The assembly drawings for the gun chamber were completed and distributed for review.  The design for the shield and internal cesiation scheme was completed, and we started detailing.  We ordered parts for a shield/mechanical movement test in vacuum.  We also ordered material to add weight into the FET system to allow for running at the higher fields expected with nitrogen implanted titanium.

Gun HVPS - The spare HVPS was crated and shipped to Glassman this week.  They will add a second multiplier stack and drive unit, test the new assembly, and ship it back to us.  We are designing the extension to the HVPS Tank and the mounting of the 2nd multiplier stack in the tank.  This effort has just started.

WBS 5 (SRF):
We received the results of the gradient measurements on the second 7-cell test cavity sent to KEK for electropolish tests.  Like the first 7-cell cavity test, this cavity also shows 20MV/m gradients after electropolishing.

Received 5 helium vessels from PHPK. Post end can award visit to Ability Engineering completed.
Completed welding of one 7 cell cavity.

WBS 6 (RF):
Zone 3 – The contract for the Cathode Power Supply is being awarded this week to Hippotronics.  No other progress was made due to parts that have not yet been received.

Zone 4 – The test software is up and running with minor corrections to be made.  The hardware punch is being worked.  With the exception of positions 7 & 8, the external coefficients have been determined and they are operating properly.  Position 8 has a circulator out for repair.  Position 7 coefficients are still being determined.

Injector 100 kW Klystron – CPI continues to have delivery problems with parts for these klystrons.  The Factory Acceptance Tests for the 1st unit have been delayed to 6/20 from 4/11.  They now plan to ship the last unit 8/15.

Injector RF – The first article 100 kW Klystron is now scheduled for factory acceptance testing on 6/20/01.  The Circulators are due 4/6/01.  The upgrade of the HVPS is still being reviewed by the vendor (PSC).  I am supplying him different design options that would be acceptable to JLab.  We may have to do this work in-house.

WBS 8 (I&C):
Progress continues to be made on the Optical Beam Position Monitors (OBPMs), time will be scheduled after the run to continue the development.

Many thanks to the software group for their work on the new trim card specification, which can be found at:  http://devweb.acc.jlab.org/controls_web/LLAPPS/Requirements/FelTrim.html.  The new system will be capable of up to 20 amps per channel; each card will have hardware strapping for full scale outputs of 1, 5, 10 or 20 amps maximum current.  This system will be purchased commercially.

Parts are on order for 40 new Ion Pump Power Supplies.  The crate design is also progressing.  Discussions continue for a redesign of the fast valve cards, this would assist the CEBAF support group since there are no spares and the programmable devices used are long since out of date.

Four (4) Drawings have been completed and submitted to Document Control.  EECAD has the Picomotor Power supply Chassis, User Lab Power Supply Chassis and the Beam Viewer Interlock Box for documentation.  Documentation continues on the Video Crosspoint Chassis wiring diagram, Alignment Mode Permit Card Schematic, and the Lab Status Lamp Control Chassis.  50 VME front panels for various active projects were shipped out for machining and silk-screening this week and are expected back in about 30 days.  Another 50 will be ordered to complete this project.

The design has begun for a new four channel VME delay/timing card.  This will generate the required signals scope triggers, sample/holds/BPMs, OBPMs.  This unit will have sub-microsecond accuracy, while a commercial unit capable of sub-nanosecond resolution will be purchased.  These two boards will be able to cover any timing needs of I&C or users.  This board will be close to the final timing card required for the new drive laser pulse controller (DLPC).  These will be located in each VME crate and driven from by fiber from the existing DLPC.  All timing and delays will be set from EPICS.

All parts have been ordered for the 8th user power supply box.  The original chassis is in the process of being documented and we should have the first set of prints early next week.  Final design for the hutch patch panel and user boxes is complete and will go to EECAD early next week.  Drawing collection continues.  We have received 8 additional drawings from Document Control this week.  Drawings for racks FL07B04 and FL07B05 will be to EECAD tomorrow.  Sixteen (16) BPMs (of 44) have been TDR'd and are in final welding.  The welded bellows for the 36 3" shielded beam viewers are also in, these were late by 3 weeks from Standard Bellows.  These are being welded on to the viewers.  Both of these fab jobs are being done by Master Machine in NN.  They are doing very well and are holding to the schedule.

The new LSS box was installed in Lab 2.  This will allow for operation of the mirror test stand.  Once lab 2 is done they will move on to lab 5 than all labs will be fully commissioned.  Vacuum design and procurements continue for the upgrade.  Drawings for the User power supply boxes are complete.  We are still waiting for some parts before we can get the new ones made.  An IOC is now working in lab 5 as a VME test stand.  This will be used to test and commission all VME hardware prior to installation in the accelerator.  It now has a fully tested GPIO card.  A computer (cast-off from AES) was reworked and commissioned as a dedicated Altera programming station, this is available for general (labwide) use and is located in the back of room 216.  There is continued progress on the GPIO card schematics (~90%) as well as other drawing for both “as-builts" and for upgrade efforts.  We received the face-plates that were sent to Service Metals for the GPIB and GPVS systems.
The rev. 3 design of the Optical Beam Position Monitor preamps is complete and is turned into EECAD. This version will be fitted into 2.5" aluminum box beam.  The Molectron Q5 detector will be used, this is
5mm round split 4 ways.  The magnification of the system will be reduced from 30 to 10.  This will create a 1mm spot for the nominal 1 cm FEL beam, the system has a spatial dynamic range of greater than one full beam diameter in any direction.  This version will be complete and ready for testing by the end of April.  Most of the parts are in for the user lab camera power supply chassis.  Design is complete for the video patch panel.  The 3" diagnostics are beginning to arrive.  The first shielded beam viewers and BPMs came in.  Discussions continue for a new design of the BPM electronics.  Initial thoughts are centered on using the AD8313 logarithmic direct down-convertor (RF to DC), a sample/hold, and a low cost DSP.  These would be located in a specifically designed compartment grouted into the new magnet stands.  These devices would then be serially linked to the IOC, preferably with a commercial network like CAN (control-area-network).

Drawings in progress:
o Camera Power Supply wiring diagram was signed off.
o Picomotor control chassis is in final review, some has been signed.
o Charge/Dump current Monitor Chassis check print has been returned to EECAD for corrections and a wiring
    diagram has been added.
o New OBPM design is in for a drawing to be made.
o Fabrication Drawings for the Ion Pump Power Supply Front Panel were signed off and sent out for fabrication.

Laser server has been up-graded.  Web work is on going.  Time has also been spent cleaning and re-stocking the I&C shop.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o DULY Research bought a new 800 MHz, large memory computer to handle the increased number of nodes of their
    magnetic models.  They are applying this computer to modeling the large Injector Dipole (GU) to obtain insight
o To this point.the GV has defied simultaneous 1 part in 10000 flatness in bulk field and field integral.
o DULY added coil leads to the drafting models.  They are workng on drawing sets for both magnets.  Any final
    adjustments to the designs from the magnet analysis will have to be propagated back through the drawing sets.
Optical Chicane Dipole (DW)
o We finished the drawing set back-checking and are starting to work on the procurement specifications.
Arc Dipoles (GY, GX, GQ)
o  AES continues progress in designing GY (180 degree dipole) finishing the coils leads area, the path length corrector
    leads and back propigating the results from magnetic modeling.
o  AES obtained perfected their 3D magnetic model of the GY with additional fine structure to the mesh to investigate
    the fields down to the 1 part in 10,000 level.  They thickened the field clamps to 3/4 inch, reducing them below
    magnetic field saturation.  They ended the month investigating field integral issues, noting that the field integral is
    decreased by the slots for the GG coils.  They have had to increase the width of the path length corrector coils to
    maintain field flatness through the ends and make the slot in the pole face (for the path length correctors wedge
    shaped to maintain a field integral that increases correctly with radius.
o  AES continues progress in designing 180 degree dipole (GY) they added more length to the faces, put a wedge
    groove into the path length corrector coil slot and widened the slot to the maximum available position as part of feed
    back from the magnetic model.  They are now laying out the leads for the path length correctors.
o  We had feed back on the draft steel specification from Bethlehem Lukens, indicating they could try to obtain our
    tight metallurgical specifications but could only guarantee slightly higher values.  They also will not be able to make
    the slabs in one heat.  The magnet steel specification was generated to purchase the 97 tons of steel and is in
    checking by Leigh Harwood.
QX (3.125" Quad)
o Measurement Probe:
o The probe was used.  It is our best quadrupole probe yet, canceling out 95% of the dipole and quadrupole
    components.  The data is in analysis.
o We continued to work on the procurement packages for this quadrupole and prepared to start pole tip chamfer
Sextupole/Octupole /Corrector
o We decided that the Octupole would most likely not be able to be generated from additional coils added to the
    sextupole because of the huge quadrupole generated.  It will most likely be a separate magnet.  Robin Wines
    continued work on the magnetic model of the Sextupole, attempting to obtain a vertical corrector dipole with the
    added coil method.  This attempt at a combination magnet was also not encouraging because of the large sextupole
    in the 2D model.  The corrector was 1 part in 100 only within half of the required width.  (A 3D model may reveal a
    compensating counter sextupole from the end fields that would extend the good field region).  In the remaining time
    that Robin has before the birth of her baby she perfected analysis of the sextupole.  We have decided to suspend
    design of the Octupole until Robin returns from maternity leave.
o We found that the magnetic model of the sextupole indicates the magnet is well within the 1% requirement at 2/3
    radius.  The reason for the improvement in analysis is the new scrip (suggested by Steve Lassiter of Physics) that
    starts analysis with the vector field potential rather than the field.
o We continued distributing diagnostics and correction elements into the lattice in consultation with David Douglas, our
    chief beam physicist.
o We started work on defining the engineering services contractor task order that will generate the design of the dipole
    vacuum chambers.
o Starting a new weekly meeting, Butch Dillon-Townes further clarified girder configurations by having all interested
    parties come together.  We started laying out the girders of the Return Leg.  These positions are the most general of
    the quad girders.  We will learn some design techniques from this exercise that allow us to design the more complex
    regions more rapidly.

WBS 10 (Wiggler):
We received the Dispersion Section for the Optical Klystron from Process Equipment Co.  It looks good.  An order to find the optimum end corrector strength for the wigglers was drafted and sent to AES.  A clogged cooling line in the dispersion section was found and fixed.  The poles are slightly askew so the jaws are being shimmed to take out the angle.

We did a preliminary magnetic measurement of the dispersion section and are analyzing the data.  A final check of the remaining drawings for the optical klystron vacuum chamber and diagnostics was completed.  Drawings should be ready for sign-off shortly.

WBS 11 (Optics):
We received the vacuum vessel for the mirror test stand, a major component for our prototype cavity mirror tests.  We have checked the vacuum vessel for the mirror test stand against the released drawings, and found only two out-of-spec dimensions, both can be tolerated.  The test pump down of the vacuum vessel late on 3/23 was successful.  We have new concepts for installing 4 mirrors in the optical cavity assemblies, and are working with the designers to see how feasible they are.  An overall system print, defining the optics subsystem regions, and positions for components, was reviewed and signed off.

We have scheduled the installation of additional utilities in User Lab 2; this will be used to support prototype cavity mirror tests.  We met with staff from Survey and Alignment to determine the placement of fiducials on the mirror mounts.  This will facilitate proper installation and setting of the cavity length.  We have opened discussions with a second vendor for the upgrade drive laser.  Draft specifications for an RFQ are being written so they can be approved and passed on to Procurement.

The bid package for two of our prototype deformable mirrors is in Procurement.  Preparations to ready User Lab 2 for cavity mirror tests is proceeding at a rapid pace; the Nd:YAG laser that will used for initial tests (a gift from AK Steel) is installed, and test run (set up to not lase).  We have new drawings and placement of components of the Upgrade optical cavity, they will be circulated for comment.  New nomenclature for some of these components was created.  Metal shields, used to protect the windows that send beam into the user labs were received from our shop, as were some parts for the manual beam shutters.  We had a service visit for the drive laser.  The vendor (Coherent) will declare the laser obsolete at the end of this month, so this was our last opportunity to have this service performed.  We now have improved performance (4W of SHG) from our well-used Nd:YLF laser rods, and a spare, new rod that produces > 7W of SHG.  We purchased spare components that should keep the laser operational for many more years.  A valve assembly for the drive laser heat exchanger began leaking building HVAC water, and will be replaced next week.

We worked with designers on design of the gimbal mounts for the optical cavity assemblies, and on the insertable mirror for the beam dump in the vault.  We received four 3" dia. uncoated, figured silicon substrates from the vendor one month earlier than promised.  Uncoated, they will be used in the Mirror Test Stand with the Nd:YAG laser to provide power loading, and to test bonding techniques required to mate the mirror to the backplane assembly.  Once coated, they will be used in the upgrade optical cavity. Preparations to ready User Lab 2 for cavity mirror tests continue.  The internal components to the mirror cassette were installed and aligned, as was the turning mirror above the cassette.  The I. & C. group continues installing cabling and LSS hardware to support use of either this laser or the FEL in the lab.


The harmonic conversion results are record-breaking (for ultrafast laser output in the visible and near-UV), and therefore deserve some additional comment.  We found that the second harmonic generation (SHG) of 1.05 micron FEL light using lithium triborate (LBO) was ~ 40% and constant up to 185 W of input power.  Maximum visible output, corrected for optics, was ~ 56W.  There was no apparent damage.  To convert to the UV, two different schemes were used.  For third harmonic generation (THG) at 350 nm, the SHG and fundamental were incident on another LBO crystal.  For fourth harmonic generation (FHG) at 262.5 nm, the SHG output was focused, then directed onto a beta barium borate (BBO) crystal.  Conversion efficiencies from the IR to the UV were about the same for THG and FHG, 9% and 8% respectively.  We briefly measured a THG output of 16 W for 100 W of input, but more often measured powers of 11W.  An IR camera showed that the second LBO crystal was heating up, and we observed a decrease in power.  This could be due to color centers being formed during irradiation; further study is needed to confirm this.  We did not have the opportunity to measure the FHG conversion under CW operation, but did use the pulsed output to irradiate H. Helvajian's samples.  We also used the 350 nm CW output to irradiate several of his samples.  As was mentioned in the Dec 8 weekly report, harmonic conversion is sensitive to the input wavelength and intensity, we made some progress in monitoring both by analyzing the FEL fundamental radiation with a spectrograph and InGaAs line scan camera (essentially a linear array) and displaying this in the control room.  A lot more work needs to be done to get this into our controls system, but the components are coming together.

It is becoming routine to measure Thomson backscattered X-rays from the FEL, both at ~5 keV and ~12 keV.  This week the first successful diffraction pattern of these X-rays off a LiF crystal was measured.  These measurements will be used to position the diffracting crystal for maximum flux for possible future IR-Xray pump-probe experiments.

We completed the winter user run on Friday March 9.  Operations the week of March 11, were concerned with difference orbit measurements (see WBS 3) and O-BPM.

Privacy and Security Notice